USE THE LAW OF EXPONENTS TO FIND THE VALUE OF "A" THAT WOULD MAKE THE FOLLOWING STATEMENT TRUE. EXPLAIN YOUR REASONING.
|X+A| = B
4^16 = 16^2X = 256^X = 4294967296
THIS QUESTION IS FROM A PRACTICE TEST FROM THE HSPA IN JERSEY CITY.
PLEASE HELP THANK YOU.

Answer by stanbon(75887)

USE THE LAW OF EXPONENTS TO FIND THE VALUE OF "A" THAT WOULD MAKE THE FOLLOWING STATEMENT TRUE. EXPLAIN YOUR REASONING.
|X+A| = B
X+A = B or X+A = -B
A = B-X or A = -(B+X)
---------------------------
4^16 = 16^2X = 256^X = 4294967296
Comment: There are too many equal signs to make any sense out of this.
Cheers,
Stan H.